If the sum of distances of a point from two perp[endicular lines in a plane is 1, the find the locus of the point.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: Line
Sol:
Let the coordinates of the point be (h, k). The equation of two perpendicular lines:
ax+by+c=0….......(1)
-bx+ay+c^{'}=0…...........(2)
Whera a, b, c & care constants. Let distance of point form (1) & (2) be d1& d2.
d_{1}+d_{2}=1
d_{1}=\frac{|ah+bk+c|}{\sqrt{a^{2}+b^{2}}}
d_{2}=\frac{|-bh+ak+c^{'}|}{\sqrt{a^{2}+b^{2}}}
\frac{|ah+bk+c|}{\sqrt{a^{2}+b^{2}}}+\frac{|-bh+ak+c^{'}|}{\sqrt{a^{2}+b^{2}}} = 1
|ah+bk+c|+|-bh+ak+c^{'}| = \sqrt{a^{2}+b^{2}}
(a-b)h+(a+b)k+c+c^{'}=\sqrt{a^{2}+b^{2}}
h\rightarrow x, b\rightarrow y
(a-b)x+(a+b)y+c+c^{'}=\sqrt{a^{2}+b^{2}}
There would be more solutions but all will be straight line.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty